### v1

# M2SE: A Multistage Multitask Instruction Tuning Strategy for Unified Sentiment and Emotion Analysis

Abstract
Sentiment analysis and emotion recognition are crucial for applications such as human-computer interaction and depression detection. Traditional unimodal methods often fail to capture the complexity of emotional expressions due to conflicting signals from different modalities. Current Multimodal Large Language Models (MLLMs) also face challenges in detecting subtle facial expressions and addressing a wide range of emotion-related tasks. To tackle these issues, we propose M2SE, a Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for general-purpose MLLMs. It employs a combined approach to train models on tasks such as multimodal sentiment analysis, emotion recognition, facial expression recognition, emotion reason inference, and emotion cause-pair extraction. We also introduce the Emotion Multitask dataset (EMT), a custom dataset that supports these five tasks. Our model, Emotion Universe (EmoVerse), is built on a basic MLLM framework without modifications, yet it achieves substantial improvements across these tasks when trained with the M2SE strategy. Extensive experiments demonstrate that EmoVerse outperforms existing methods, achieving state-of-the-art results in sentiment and emotion tasks. These results highlight the effectiveness of M2SE in enhancing multimodal emotion perception. The dataset and code are available at https://github.com/xiaoyaoxinyi/M2SE.
Index Terms:
Multimodal Large Language models, Multitask, Sentiment, Emotion.
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Sentiment analysis and emotion recognition are critical components in a variety of domains, including human-computer interaction[1], depression detection[2], and more. These tasks involve identifying sentiments and emotional states in human communication. While unimodal approaches—those that rely on individual modalities such as face[3], text[4], or audio[5]—have achieved notable success in sentiment and emotion detection, they are inherently limited in their ability to fully capture the complexity of human emotions. Each modality alone provides only a partial view of the emotional landscape, making it difficult to analyze data in a comprehensive and holistic manner. Furthermore, unimodal systems are prone to misinterpretations, especially in cases where emotions are expressed indirectly or ambiguously, such as with sarcasm[6], irony, or subtle emotional cues. These challenges highlight the need for more integrated and multimodal approaches to sentiment and emotion analysis that can overcome the limitations of single-modality methods.
Existing methods for fusing multimodal information typically rely on cross-modal interactions to learn from different modalities and ultimately perform sentiment analysis or emotion recognition tasks[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. However, there is an inherent relationship between sentiment and emotion that these methods often struggle to capture. Simply completing sentiment or emotion tasks in isolation does not effectively model the complex interplay between specific emotions and corresponding sentiments.
Moreover, while current Multimodal Large Language Models (MLLMs) demonstrate strong performance in tasks such as visual language understanding [30, 31, 32], they often underperform in tasks like multimodal sentiment analysis and emotion recognition in zero-shot settings. As a result, fine-tuning remains essential. As shown in Figure 1, although existing MLLMs achieve promising results when fine-tuned for a single task, their performance on other tasks frequently deteriorates. This underscores the need for a general-purpose MLLM capable of performing well across a variety of sentiment- and emotion-related tasks, and capturing both sentiment and emotion in a more integrated and comprehensive manner.
Figure 1: A comparison of MLLMs trained solely on the MSA task, solely on the ER task, and using the M2SE strategy. (a) In the MSA task, the model trained on MSA performs well, while the model trained on ER performs poorly. (b) In the ER task, the model trained on ER performs well, while the model trained on MSA performs poorly. (c) EmoVerse, trained with the M2SE strategy, not only excels on both MSA and ER tasks but also successfully performs the emotion reason inference task.
To fine-tune any MLLM with multitask capabilities, we explore the relationships between different tasks and propose a novel strategy called M2SE: Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for common MLLMs. Through our exploration of task relationships, M2SE applies a multistage approach, where different tasks are introduced at each stage of training, and each task is assigned a different sampling rate based on its importance. By using this strategy, we train Emotion Universe (EmoVerse), a model based on traditional MLLM architecture that integrates both visual and textual modalities. Notably, EmoVerse achieves strong performance on sentiment and emotion tasks without requiring any specialized multitask learning module. Instead, it leverages the basic MLLM structure and the M2SE strategy to effectively analyze and reason about complex emotional contexts.
In addition, existing datasets often fail to provide the comprehensive and nuanced data necessary for training models that can effectively handle multiple tasks. This underscores the urgent need for a more robust dataset capable of supporting multitask learning in sentiment and emotion analysis. To address this gap, we introduce the Emotion Multitask (EMT) dataset, specifically designed to support MLLMs across five distinct tasks: multimodal sentiment analysis (MSA), emotion recognition (ER), facial expression recognition (FER), emotion reasoning inference (ERI), and emotion cause-pair extraction (ECPE) [33, 34]. By incorporating multiple tasks, the EMT dataset facilitates the development of models that can recognize, reason, and infer the causes of sentiment and emotion.
In summary, our main contributions are as follows:
  • We construct the EMT dataset. Each piece of data in this dataset contains queries for five tasks and the corresponding labels. The EMT dataset encourages the model to capture the relationship between sentiment and emotion from the perspective of different tasks, enabling the model to extract richer contextual information.
  • We introduce the M2SE strategy, which enables effective multitask learning by leveraging multistage training on diverse emotion and sentiment-related tasks. To the best of our knowledge, this is the first time the ECPE task has been introduced within the emotion domain as part of a multitask learning framework.
  • We develop the EmoVerse model. EmoVerse unifies tasks in the sentiment and emotion domains by employing the M2SE strategy. To ensure better adaptability to edge devices, we train the model with two different parameter configurations: 4B and 8B.
  • EmoVerse demonstrates outstanding performance across a wide range of tasks. On the CMU-MOSEI dataset for MSA[35], it achieves an Acc2 score of 88.51%. Similarly, on the MELD dataset for ER[36], it attains a weighted F1 score of 66.74. Additionally, ablation studies further validate the effectiveness of the M2SE strategy.
Figure 2: An example from the EMT dataset, which includes five tasks: sentiment analysis, emotion recognition, emotion reason inference, facial expression recognition, and emotion cause-pair extraction———along with its data construction process.
II Related Work
II-A Multimodal Large Language Models (MLLMs)
Multimodal Large Language Models (MLLMs) have gained significant attention due to their powerful reasoning capabilities. These models are typically trained for general-purpose tasks using a standard framework that includes modal encoders, a large language model (LLM), and a mapping layer (e.g., Linear [30], MLP [32, 37], or Q-Former [38]). While they excel in general tasks, MLLMs struggle with MSA and ER due to insufficient training on sentiment-specific datasets and emotion-related knowledge. Recent efforts have focused on improving MLLMs by incorporating multimodal emotion datasets and emotion reasoning tasks [39, 40]. However, these models still face challenges in leveraging contextual information effectively, particularly in sentiment analysis, where understanding the emotional nuances is crucial.
II-B Multitask Learning (MTL)
Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others [41]. In affective computing, where multimodal inputs (e.g., text, audio, and visual frames) carry diverse and distinct information, MTL frameworks aim to leverage the synergies between tasks for improved performance. For instance, the MTMM-ES framework [42] jointly performs sentiment and emotion analysis, showing improved results over single-task models. UniMSE [11] unifies MSA and ER within a single framework. However, it relies solely on labeled data to establish the relationship between sentiment and emotion, without fully utilizing multimodal connections. M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture. Recently, Emotion-LLaMA [39] has shown promising results in ERC by incorporating multitask learning, including emotion cause inference. However, despite its success in ERC, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
Figure 3: Overview of the EmoVerse training process and model architecture. For different training stages, we assign tasks to the data based on their respective sampling rates, which are indicated in parentheses. The EmoVerse model is then trained using the assigned data. The model architecture include visual encoder, linear projector, and LLM.
III Methodology
This section presents our work, which consists of three key components: the construction of the EMT dataset, the formulation of the M2SE strategy, and the development of the EmoVerse model.
III-A EMT datasets
The EMT dataset is constructed by integrating data from the CMU-MOSEI, MELD, and ECF2.0[33] datasets to support five distinct tasks: MSA, ER, FER, ERI, and ECPC. Figure 2 illustrates a sample from the EMT dataset.
For the MSA and ER tasks, we use the original sentiment labels and emotion annotations from the source datasets as task labels. Due to the complexity of annotating the ECPE task, we refer to the annotation method of the ECF dataset and manually annotate a portion of the data. This is then combined with the existing ECF2.0 data to form the ECPE portion of the EMT dataset for further emotional causality analysis.
In the FER task, we adapt methods from the MERR dataset[39] by extracting peak frames based on the Action Units (AUs) calculated for each frame using the OpenFace tool. The composite score for each frame is determined by summing the AU values, as shown in the following equation:
$S_{f}=\sum_{i=1}^{n}\text{AU}_{i}(f)$ (1)
where $S_{f}$ is the composite score of frame $f$, and $\text{AU}_{i}(f)$ represents the Action Unit value for the $i$-th facial action unit at frame $f$. The peak frame with the highest score is selected to represent the most expressive moment. For videos featuring multiple characters, such as those in the MELD dataset, we improve the method by extracting peak frames separately for each character and then selecting the final peak frame by comparing the highest composite scores from all characters:
$S_{f_{\text{final}}}=\max\left(\{S_{f}^{(1)},S_{f}^{(2)},\dots,S_{f}^{(k)}\}\right)$ (2)
where $S_{f}^{(i)}$ is the composite score for character $i$ and $k$ represents the total number of characters in the frame. The final peak frame is the one with the highest composite score among all characters.
After selecting the peak frames, we extract the facial expression captions by combining the AUs from the peak frame with the predefined AU list corresponding to the emotion. Let $\text{AU}_{\text{frame}}(f)$ represent the set of AUs from the peak frame for a given emotion, and $\text{AU}_{\text{emotion}}(e)$ represent the predefined AU list for a specific emotion $e$. The common AUs are obtained by taking the intersection of the two sets:
$\text{AU}_{\text{common}}(f,e)=\text{AU}_{\text{frame}}(f)\cap\text{AU}_{\text{emotion}}(e)$ (3)
where $\text{AU}_{\text{common}}(f,e)$ is the set of AUs common to both the peak frame and the emotion-specific AU list. These common AUs are then mapped to textual information to generate the FER annotations for the task. This approach enables the model to link the facial expressions captured in the peak frames with the corresponding emotional state.
In the ERI task, we input the video into the Internvl2-8B  [^1] https://huggingface.co/OpenGVLab/InternVL2-8B model to generate detailed descriptions of both the characters and the scene. These descriptions allow the model to analyze emotional causes by considering facial expressions, speech, and environmental context. Finally, we use the LLaMA 3.1-8B  [^2] https://github.com/meta-llama/llama3 model to infer the underlying causes of the characters’ emotions by synthesizing the character and scene descriptions.
We perform a thorough preprocessing and quality control process on the raw data from the three datasets, systematically eliminating problematic or noisy data. Simultaneously, we conduct a manual curation of high-quality samples, following the aforementioned methodology. This involves refining the dataset to ensure accuracy and consistency, while extracting the necessary labels for FER and ERI. The final dataset, after this meticulous refinement, is outlined in Table I, including the number of tasks per category. Note that each data sample can be associated with multiple tasks. Table I below breaks down the number of tasks that can be used.
Additionally, we transform the dataset into a structured format, which includes three key elements: query, response, and images/videos. This format facilitates the instruction tuning of MLLMs.
TABLE I: The Number of Entries in EMT Dataset
Tasks Number of Entries Sentiment Analysis 25,859 Emotion Recognition 25,859 Facial Expression Recognition 15,870 Emotion Reason Inference 4,839 Emotion Cause-Pair Extraction 7,081 Total (Summed) 32,940
III-B EmoVerse
To validate the effectiveness of the M2SE strategy, we select the most widely used MLLM framework without introducing any multitask modifications to its architecture. This ensures that any observed performance improvements can be directly attributed to the proposed training strategy, rather than architectural changes. The proposed MLLM architecture, EmoVerse, as shown in Figure 3, consists of a visual encoder, a linear projector, and the LLM.
We train EmoVerse following M2SE strategy using the EMT datasets. For a data sample $D$, the input consists of the pair $\{v,t\}$, where $v$ represents the video and $t$ represents the text. First, the video $v$ is processed to extract visual tokens using a Vision Transformer (ViT)[44], producing a set of visual tokens $T_{v}$. These visual tokens $T_{v}$ are then passed through a linear transformation $\mathbf{W}_{v}$ to align the visual features with the feature space of the LLM. Mathematically, this can be expressed as:
$T^{\prime}_{v}=\mathbf{W}_{v}T_{v}$ (4)
where $T^{\prime}_{v}$ represents the aligned visual features.
Next, the text $t$, task identifiers, and instructions are tokenized and passed through the embedding layer of the LLM, producing the tokenized representation $T_{t}$. The tokenized visual features $T^{\prime}_{v}$ and the tokenized text $T_{t}$ are then concatenated to form the final input to the model:
$\mathbf{X}=[T^{\prime}_{v};T_{t}]$ (5)
where $\mathbf{X}$ is the combined input. The model processes this concatenated input and outputs the response $\hat{y}$, which corresponds to the task at hand:
$\hat{y}=\mathcal{M}(\mathbf{X})$ (6)
where $\mathcal{M}$ is the model and $\hat{y}$ is the model’s prediction or answer.
III-C Multi-Stage Multitask Sentiment and Emotion Instruction Tuning
III-C1 Stage 1: Multitask Pretraining
In the first phase of training, the model focuses on three key tasks: MSA, ER, and FER. We randomly select 15,000 samples from the EMT dataset for this phase, giving higher sampling rates to the MSA and FER tasks, while assigning a lower sampling rate to the ER task. The primary objective of this stage is to enable the model to recognize facial expressions and, through these expressions, capture the relationship between sentiment and emotion. Specifically, the FER task helps the model learn fine-grained facial features, which in turn improves performance in both the MSA and ER tasks by reinforcing the connection between emotional cues and sentiment.
III-C2 Stage 2: Multitask Reason Fine-Tuning
Following the first phase, where the model learns the fundamental relationship between sentiment and emotion as well as facial expression recognition, the second phase introduces four primary tasks: MSA, ER, ERI, and ECPE. For this phase, we select the remaining samples from the EMT dataset, randomly distributing them across the tasks. Notably, the MSA task is assigned a very small sampling rate, while the other three tasks—ER, ERI, and ECPE—are given higher and equal sampling rates. The ERI task enhances the model’s ability to reason about emotional causes by analyzing speech, facial expressions, and contextual information, building on the facial expression recognition learned in Stage 1. On the other hand, the ECPE task provides deeper insights into the causes of emotions by linking emotional responses with their underlying causes in conversations. Through the extraction of emotion cause-pairs, ECPE fosters better integration of multimodal and contextual information, helping the model understand the complex relationships between sentiment, emotion, and context. The details of Stage 1 and Stage 2 can be found in the Table II.
TABLE II: Stage 1 and Stage 2 with Tasks and Task Identifiers
Different Tasks Stages Task Identifier Stage 1 Stage 2 Sentiment Analysis <sentiment> Emotion Recognition <emotion> Facial Expression Recognition × <caption> Emotion Reason Inference × <reason> Emotion Cause-Pair Extraction × <emotion cause-pair>
IV Experiments
IV-A Experiments Setup
EmoVerse is trained solely on the EMT dataset without the addition of any external data, and it is tested on the following task-specific test sets.
IV-A1 Sentiment Analysis.
We use the CMU-MOSEI datasets for MSA. The CMU-MOSEI dataset is an extended version of the CMU-MOSI dataset. It encompasses 3228 videos from 1000 speakers, comprising a total of 23453 sentences. Similar to the CMU-MOSI dataset, it covers multiple emotional dimensions, employing the same annotation labeling approach. The results for Acc-2 calculated based on negative/non-negative (N/N) scheme, while the results on the right are calculated according to negative/positive (N/P) definition.
IV-A2 Emotion Recognition.
We use the MELD datasets for ER. The MELD dataset is an emotional dialogue dataset comprising 1433 dialogue segments and 13708 utterances from movies and TV shows. Each statement in the dialogue is labeled with one of seven emotions: anger, disgust, sadness, joy, neutral, surprise, and fear. Furthermore, the MELD data set provides annotations for emotional polarity (positive, negative, and neutral) for each utterance. We use Acc and weighted f1-sorce for evaluation.
IV-A3 Emotion Cause-pair Extraction.
We use the ECF2.0 datasets for ECPE. The ECF2.0 dataset is a comprehensive resource for emotion cause-pair extraction, comprising 1,715 conversations and 16,720 utterances. Each utterance in the dataset is annotated with emotion cause-pair information, which helps in understanding the emotional context and its causes. The dataset is divided into training and evaluation sets, with 1,374 conversations and 13,619 utterances in the training set, and 341 conversations and 3,101 utterances in the test set. We refer to previous evaluations and assess the ECF2.0 test set using the F1 Score and Weighted F1 Score.
IV-B Implementation Details
For the visual encoder, we utilize pre-train ViT with input images resized to $448\times 448$ pixels. During the fine-tuning process, the visual backbone is frozen, with the focus placed on training the linear projection layer. For LLM, we employ Phi-3-Mini[45] for EmoVerse-4B and InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLMfor EmoVerse-8B, both augmented with LoRA[46] for parameter-efficient fine-tuning.
In each stage of training, we train for 2 epochs with LoRA hyperparameters set to $r=8$ and $a=32$. The learning rate is set to $1\times 10^{-5}$, and we use a warm-up cosine schedule for the learning rate decay. For EmoVerse-4B, the training is conducted on 2 NVIDIA RTX 3090 GPUs, requiring approximately 48 hours. For EmoVerse-8B, the training is executed on a NVIDIA A100 GPU, taking around 55 hours.
IV-C Results
IV-C1 Performance comparison.
For MSA task, we compare our EmoVerse model with several SOTA baselines, including TFN[7], MulT[12], PMR[21], MISA[17], MMIM[14], Self-MM [13], BBFM[15], FDMER[16],CONKI[18], DMD[20], ConFEDE[19], and MRC-D3AE[27]. In ER task, we evaluate our model against baseline models with DialogueGCN [22], DialogueCRN[23], DAG-ERC[8], MM-DFN[24], EmoCaps[9], ERMC-DisGCN[25], MMGCN[26], GA2MIF[10], UniMSE[11], and FacialMMT[47].
Additionally, for both MSA and ER, we select advanced models capable of processing both images and videos. These include LLaVA-Video [48] and InternVL2 [49], both of which are designed to handle multimodal input efficiently. The performance comparison results, presented in Table III and Table IV, clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, surpassing all the aforementioned baseline models in terms of accuracy and robustness.
For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task. The Table V shows the results of ECPE.
TABLE III: The performance comparison on CMU-MOSEI dataset.
Model Acc2 TFN - / 82.5 MulT - / 82.5 Self-MM 82.81 / 85.17 MMIM 82.24 / 85.97 BBFN - / 86.20 FDMER - / 86.10 MISA 83.60 / 85.50 ConKI 82.73 / 86.25 ConFEDE 81.65 / 85.82 DMD - / 86.60 PMR - / 83.30 MRC-D3AE 83.10 / 85.50 LLaVA-Video - / 82.82 InternVL2-4B - / 82.09 InternVL2-8B - / 85.30 EmoVerse-4B 85.51 / 87.80 EmoVerse-8B 85.93 / 88.51
TABLE IV: The performance comparison on MELD dataset.
Model Acc W.F1-Score DialogueGCN 59.46 58.10 DialogueCRN 60.73 58.39 DAG-ERC - 63.65 MM-DFN - 59.46 EmoCaps - 64.00 ERMC-DisGCN - 64.22 MMGCN 62.49 58.65 GA2MIF - 58.94 UniMSE 65.09 65.51 FacialMMT - 66.58 LLaVA-Video 23.87 22.65 InternVL2-4B 39.46 40.59 InternVL2-8B 42.18 44.13 EmoVerse-4B 67.28 66.58 EmoVerse-8B 67.78 66.74
TABLE V: The performance comparison on ECF2 dataset.
Model W.F1-Score F1-Score LLaVA-Video 22.59 22.78 InternVL2-4B 33.24 33.35 InternVL2-8B 40.57 40.51 EmoVerse-4B 72.30 72.21 EmoVerse-8B 73.54 73.62
IV-C2 Analysis of Emotion Reasoning.
To demonstrate the qualitative performance of EmoVerse, we present a detailed comparison of emotion reasoning results across various models. Figure 4 showcases the emotion reasoning outputs of five models: GPT-4o, Video-LLaMA2[50], LLaMA-Video, InternVL2, and EmoVerse. The video depicts a woman displaying a strong reaction of surprise, highlighting the models’ ability to infer emotional states from multimodal inputs.
GPT-4o and Video-LLaMA2 both classify the emotion in the video as ”confident” or ”assertive.” However, LLaMA-Video and InternVL2 correctly identify the emotion as ”surprise.” Despite this, these two models fail to fully integrate multimodal information for ERI. For instance, LLaMA-Video makes its inference solely based on facial expressions, while InternVL2 briefly acknowledges the facial expressions without providing a detailed explanation of the emotional cause, instead relying solely on textual cues for emotion reasoning. In contrast, EmoVerse effectively combines both multimodal video and text information to infer the cause of the emotion, yielding a correct and comprehensive emotional interpretation.
Figure 4: Comparison of different models on ERI. It can be seen that both GPT-4o and Video-LLaMA2 make errors in ER. While Llava-Video and InternVL2 correctly classify the emotions, their descriptions of the reasons behind the emotions are incomplete. In contrast, EmoVerse not only correctly classifies the emotions but also provides a comprehensive description of the reasons behind them.
IV-D Ablation Evaluation
We use EmoVerse-4B to do the following ablation experiments.
IV-D1 Comparison of the M2SE Strategy with Single-Task Fine-Tuning.
To evaluate the impact of the M2SE strategy within the same model framework and compare its performance across various tasks, we conduct an ablation study. As shown in Table VI, the tasks in parentheses represent the fine-tuning tasks, with each task corresponding to the same datasets used in our evaluation. ”Z.S.” stands for zero-shot. The M2SE strategy is trained using the EMT dataset. The results demonstrate that, although the performance on some metrics may not exceed that of single-task learning, EmoVerse trained with the M2SE strategy consistently performs better overall compared to other models.
TABLE VI: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21
IV-D2 Effect of the FER Task.
To explore the contribution of the FER task to MSA, we conduct an ablation experiment to compare the effects of increasing and decreasing the FER sampling ratio on MSA performance, using the same training data volume in the first stage. For a low sampling ratio, we denote it as ”lf,” and for a high sampling ratio, we denote it as ”hf.” As shown in Table VII, under the same data volume, increasing the FER task ratio—despite the model performing fewer MSA tasks—still leads to stronger learning performance. This indicates that even with fewer MSA samples, the contribution of FER significantly enhances the model’s ability to perform sentiment analysis.
TABLE VII: Effect of the FER Task.
Model Acc2 EmoVerse (lf) 77.09 / 87.47 EmoVerse (hf) 82.61 / 87.50
IV-D3 Effect of the Choice of Tasks in the Second Stage.
To investigate how task configurations impact model performance and better understand cross-task interactions, we conduct an ablation study in the second stage with three setups: ”t1” (only ER task), ”t3” (ER, ERI, and ECPE tasks), and ”t4” (all four tasks). Results, as shown in Table VIII, show that focusing solely on ER (t1) improves ER performance but weakens MSA and emotion reasoning tasks. Adding ERI and ECPE (t3) slightly lowers MSA and ER performance but enhances the model’s emotion reasoning ability. When MSA is reintroduced with a small sampling rate (t4), the model effectively integrates cross-task knowledge, improving performance across all tasks and surpassing first-stage MSA results. This demonstrates the importance of balanced task integration for holistic model improvement.
TABLE VIII: Effect of the Distribution of Tasks on EmoVerse.
Model CMU-MOSEI MELD ECF2.0
Acc2 Acc W.F1 F1 W.F1
EmoVerse (f) 82.61 / 87.50 - - - -
EmoVerse (t1) 81.68 / 87.20 66.74 66.24 25.40 26.69
EmoVerse (t3) 79.0 / 86.78 66.97 65.67 69.85 69.95
EmoVerse (t4) 85.51 / 87.80 67.78 66.74 72.30 72.21
V Conclusion
In this paper, we propose the M2SE strategy to guide the training process, enabling the model to learn from multiple sentiment- and emotion-related tasks across different stages. To support this, we construct the EMT dataset, which includes a large number of annotated samples spanning a wide range of tasks. Additionally, we introduce EmoVerse, a unified MLLM designed to enhance emotion perception by effectively integrating both visual and textual data. By leveraging multitask and multistage learning, EmoVerse unifies sentiment and emotion tasks, achieving superior performance in MSA, ER, and EPCE. It outperforms existing models on benchmark datasets such as CMU-MOSEI and MELD. The experimental results highlight the effectiveness of the M2SE strategy in enabling the model to capture and reason about complex emotional contexts, showcasing its potential to drive advancements in emotion-aware AI applications.
Acknowledgments
This work was supported by the Key R&D Program of Shandong Province, China (Major Scientific and Technological Innovation Project) (NO.2022CXGC010504) and National Natural Science Foundation of China under Grant 61301253.

### v2

@@
 Multimodal Large Language models, Multitask, Sentiment, Emotion.
-publicationid: pubid: $0$–$00$ 2024 IEEE
 I Introduction
@@
 M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture.
-Recently, Emotion-LLaMA [39] has shown promising results in ERC by incorporating multitask learning, including emotion cause inference.
-However, despite its success in ERC, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
+Recently, Emotion-LLaMA [39] has shown promising results in MER by incorporating multitask learning, including emotion reason inference.
+However, despite its success in MER, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
 Figure 3: Overview of the EmoVerse training process and model architecture.
@@
 TABLE VI: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
-Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21
+Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.28 66.58 72.30 72.21
 IV-D2 Effect of the FER Task.

### v3

@@
-M2SE: A Multistage Multitask Instruction Tuning Strategy for Unified Sentiment and Emotion Analysis
+for Sentiment and Emotion Understanding
 Abstract
-Sentiment analysis and emotion recognition are crucial for applications such as human-computer interaction and depression detection.
-Traditional unimodal methods often fail to capture the complexity of emotional expressions due to conflicting signals from different modalities.
-Current Multimodal Large Language Models (MLLMs) also face challenges in detecting subtle facial expressions and addressing a wide range of emotion-related tasks.
-To tackle these issues, we propose M2SE, a Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for general-purpose MLLMs.
-It employs a combined approach to train models on tasks such as multimodal sentiment analysis, emotion recognition, facial expression recognition, emotion reason inference, and emotion cause-pair extraction.
-We also introduce the Emotion Multitask dataset (EMT), a custom dataset that supports these five tasks.
-Our model, Emotion Universe (EmoVerse), is built on a basic MLLM framework without modifications, yet it achieves substantial improvements across these tasks when trained with the M2SE strategy.
-Extensive experiments demonstrate that EmoVerse outperforms existing methods, achieving state-of-the-art results in sentiment and emotion tasks.
-These results highlight the effectiveness of M2SE in enhancing multimodal emotion perception.
-The dataset and code are available at https://github.com/xiaoyaoxinyi/M2SE.
-Index Terms:
-Multimodal Large Language models, Multitask, Sentiment, Emotion.
-I Introduction
-Sentiment analysis and emotion recognition are critical components in a variety of domains, including human-computer interaction[1], depression detection[2], and more.
-These tasks involve identifying sentiments and emotional states in human communication.
-While unimodal approaches—those that rely on individual modalities such as face[3], text[4], or audio[5]—have achieved notable success in sentiment and emotion detection, they are inherently limited in their ability to fully capture the complexity of human emotions.
-Each modality alone provides only a partial view of the emotional landscape, making it difficult to analyze data in a comprehensive and holistic manner.
-Furthermore, unimodal systems are prone to misinterpretations, especially in cases where emotions are expressed indirectly or ambiguously, such as with sarcasm[6], irony, or subtle emotional cues.
-These challenges highlight the need for more integrated and multimodal approaches to sentiment and emotion analysis that can overcome the limitations of single-modality methods.
-Existing methods for fusing multimodal information typically rely on cross-modal interactions to learn from different modalities and ultimately perform sentiment analysis or emotion recognition tasks[7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
-However, there is an inherent relationship between sentiment and emotion that these methods often struggle to capture.
-Simply completing sentiment or emotion tasks in isolation does not effectively model the complex interplay between specific emotions and corresponding sentiments.
-Moreover, while current Multimodal Large Language Models (MLLMs) demonstrate strong performance in tasks such as visual language understanding [30, 31, 32], they often underperform in tasks like multimodal sentiment analysis and emotion recognition in zero-shot settings.
-As a result, fine-tuning remains essential.
-As shown in Figure 1, although existing MLLMs achieve promising results when fine-tuned for a single task, their performance on other tasks frequently deteriorates.
-This underscores the need for a general-purpose MLLM capable of performing well across a variety of sentiment- and emotion-related tasks, and capturing both sentiment and emotion in a more integrated and comprehensive manner.
-Figure 1: A comparison of MLLMs trained solely on the MSA task, solely on the ER task, and using the M2SE strategy.
-(a) In the MSA task, the model trained on MSA performs well, while the model trained on ER performs poorly.
-(b) In the ER task, the model trained on ER performs well, while the model trained on MSA performs poorly.
-(c) EmoVerse, trained with the M2SE strategy, not only excels on both MSA and ER tasks but also successfully performs the emotion reason inference task.
-To fine-tune any MLLM with multitask capabilities, we explore the relationships between different tasks and propose a novel strategy called M2SE: Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy for common MLLMs.
-Through our exploration of task relationships, M2SE applies a multistage approach, where different tasks are introduced at each stage of training, and each task is assigned a different sampling rate based on its importance.
-By using this strategy, we train Emotion Universe (EmoVerse), a model based on traditional MLLM architecture that integrates both visual and textual modalities.
-Notably, EmoVerse achieves strong performance on sentiment and emotion tasks without requiring any specialized multitask learning module.
-Instead, it leverages the basic MLLM structure and the M2SE strategy to effectively analyze and reason about complex emotional contexts.
-In addition, existing datasets often fail to provide the comprehensive and nuanced data necessary for training models that can effectively handle multiple tasks.
-This underscores the urgent need for a more robust dataset capable of supporting multitask learning in sentiment and emotion analysis.
-To address this gap, we introduce the Emotion Multitask (EMT) dataset, specifically designed to support MLLMs across five distinct tasks: multimodal sentiment analysis (MSA), emotion recognition (ER), facial expression recognition (FER), emotion reasoning inference (ERI), and emotion cause-pair extraction (ECPE) [33, 34].
-By incorporating multiple tasks, the EMT dataset facilitates the development of models that can recognize, reason, and infer the causes of sentiment and emotion.
+Sentiment and emotion understanding are essential to applications such as human-computer interaction and depression detection.
+While Multimodal Large Language Models (MLLMs) demonstrate robust general capabilities, they face considerable challenges in the field of affective computing, particularly in detecting subtle facial expressions and handling complex emotion-related tasks, such as emotion reason inference and understanding emotions in long-context scenarios.
+Furthermore, there is a lack of a unified MLLM that can effectively handle both sentiment and emotion-related tasks.
+To address these challenges, we explore multi-task training strategies for MLLMs in affective computing and introduce Emotion Universe (EmoVerse), an MLLM designed to handle a broad spectrum of sentiment and emotion-related tasks.
+In addition, EmoVerse is capable of deeply analyzing the underlying causes of emotional states.
+We also introduce the Affective Multitask (AMT) Dataset, which supports multimodal sentiment analysis, multimodal emotion recognition, facial expression recognition, emotion reason inference, and emotion cause-pair extraction tasks.
+Extensive experiments demonstrate that EmoVerse outperforms existing methods, achieving state-of-the-art results in sentiment and emotion-related tasks.
+The code is available at https://github.com/liaolea/EmoVerse.
+00footnotetext: $\dagger$ Equal contribution. * Corresponding authors.
+1 Introduction
+Sentiment and emotion understanding are crucial for applications such as human-computer interaction and mental health monitoring.
+Traditional unimodal approaches, which rely solely on facial expressions Jiang et al.
+(2020), text Lei et al.
+(2023), or audio Hsu et al.
+(2021), have made progress but remain inherently limited, as each modality captures only a partial view of human emotions, making it challenging to achieve a comprehensive understanding.
+To address these limitations, researchers Yang et al.
+(2023); Li et al.
+(2023c); Lv et al.
+(2021); Ghosal et al.
+(2019); Hu et al.
+(2021a, 2022a); Sun et al.
+(2021); Hu et al.
+(2021c); Zhang et al.
+(2024a); Cheng et al.
+(2023); Wang et al.
+(2024b) have increasingly turned to multimodal methods that integrate multiple modalities for sentiment analysis or emotion recognition.
+While these approaches have enhanced classification accuracy, they predominantly focus on identifying sentiment or emotions, rather than delving into their underlying causes or reasoning about emotional contexts.
+Recent multimodal large language models (MLLMs) have made significant strides in general visual-language understanding Liu et al.
+(2024).
+However, in the field of affective computing, their zero-shot performance remains limited, particularly in tasks such as facial expression recognition, emotion reason inference, and contextual emotion understanding.
+Moreover, there is currently a lack of an MLLM capable of performing multiple emotion-related tasks.
+MLLMs can effectively perform multiple visual tasks in a zero-shot setting after visual instruction fine-tuning.
+However, in the field of affective computing, existing MLLMs continue to struggle with sentiment understanding after fine-tuning with emotion-related instruction data, as shown in Figure 1 and exemplified by models such as Emotion-LLaMA Cheng et al.
+(2024a).
+This discrepancy highlights a broader issue: existing MLLMs fail to capture the intricate relationship between sentiment and emotion.
+A promising solution to this challenge is multitask learning, where the model is trained simultaneously on multimodal sentiment analysis (MSA) and multimodal emotion recognition (MER) tasks.
+An example of this approach is UniMSE Hu et al.
+(2022b), which demonstrates the effectiveness of multitask learning.
+However, a significant open challenge remains: how to optimally leverage multiple tasks to enhance MLLM training?
+Figure 1: Comparison of MLLMs trained on the MSA task, MER task, and using the M2SE strategy.
+(a) The MSA-trained model performs well on MSA but poorly on MER.
+(b) The ER-trained model performs well on MER but poorly on MSA.
+(c) EmoVerse, trained with the M2SE strategy, excels on both MSA and MER tasks and successfully performs emotion reasoning inference.
+In response, we propose to train a general-purpose MLLM for affective computing using multitask learning.
+In addition to MSA and MER, we incorporate facial expression recognition (FER), emotion reasoning inference (ERI), and emotion cause-pair extraction (ECPE) tasks, aiming to improve the model’s capabilities in facial expression analysis, causal reasoning, and long-context comprehension.
+The ECPE task requires the model to identify the emotion in a specific sentence within a dialogue and determine the cause of that emotion, which is linked to a previous sentence with multimodal setting.
+This task demands a high level of reasoning and context understanding from the model.
+By examining the relationships among these tasks, we propose an optimal training strategy called Multistage Multitask Sentiment and Emotion Instruction Tuning Strategy (M2SE), which is designed to enable MLLMs to achieve strong performance across a variety of affective tasks.
+This training strategy is adaptable to different MLLMs.
+Based on the M2SE strategy, we develop a MLLM named Emotion Universe (EmoVerse).
+Additionally, existing datasets are often insufficient for training MLLMs in multitask settings due to their lack of diversity and task coverage.
+To address this issue, we construct the Affective Multitask (AMT) dataset, which includes the five tasks mentioned above.
+By incorporating these tasks, the AMT dataset facilitates the development of models capable of recognizing, reasoning, and inferring the causes of sentiment and emotion.
 In summary, our main contributions are as follows:
-  • We construct the EMT dataset.
+  • We construct the AMT dataset.
 Each piece of data in this dataset contains queries for five tasks and the corresponding labels.
-The EMT dataset encourages the model to capture the relationship between sentiment and emotion from the perspective of different tasks, enabling the model to extract richer contextual information.
-  • We introduce the M2SE strategy, which enables effective multitask learning by leveraging multistage training on diverse emotion and sentiment-related tasks.
-To the best of our knowledge, this is the first time the ECPE task has been introduced within the emotion domain as part of a multitask learning framework.
-  • We develop the EmoVerse model.
-EmoVerse unifies tasks in the sentiment and emotion domains by employing the M2SE strategy.
-To ensure better adaptability to edge devices, we train the model with two different parameter configurations: 4B and 8B.
+The AMT dataset encourages the model to capture the relationship between sentiment and emotion from the perspective of different tasks, enabling the model to extract richer contextual information.
+  • We explore the relationships between different sentiment and emotion-related tasks within a multitask learning framework.
+Through this exploration, we identify the optimal strategy, called M2SE.
+This exploration is the first multitask investigation in the field of affective computing.
+  • We develop the EmoVerse model. EmoVerse unifies tasks in the sentiment and emotion domains by leveraging the M2SE strategy.
   • EmoVerse demonstrates outstanding performance across a wide range of tasks.
-On the CMU-MOSEI dataset for MSA[35], it achieves an Acc2 score of 88.51%.
-Similarly, on the MELD dataset for ER[36], it attains a weighted F1 score of 66.74.
-Additionally, ablation studies further validate the effectiveness of the M2SE strategy.
-Figure 2: An example from the EMT dataset, which includes five tasks: sentiment analysis, emotion recognition, emotion reason inference, facial expression recognition, and emotion cause-pair extraction———along with its data construction process.
-II Related Work
-II-A Multimodal Large Language Models (MLLMs)
+On the CMU-MOSEI dataset for MSAZadeh et al.
+(2018), it achieves an Acc2 score of 88.51%.
+Similarly, on the MELD dataset for MERPoria et al.
+(2018), it attains a weighted F1 score of 66.74.
+Figure 2: The construction process of the AMT dataset.
+The AMT dataset includes five tasks: multimodal sentiment analysis, multimodal emotion recognition, emotion reason inference, facial expression recognition, and emotion cause-pair extraction.
+2 Related Work
+2.1 Multimodal Large Language Models
 Multimodal Large Language Models (MLLMs) have gained significant attention due to their powerful reasoning capabilities.
-These models are typically trained for general-purpose tasks using a standard framework that includes modal encoders, a large language model (LLM), and a mapping layer (e.g., Linear [30], MLP [32, 37], or Q-Former [38]).
-While they excel in general tasks, MLLMs struggle with MSA and ER due to insufficient training on sentiment-specific datasets and emotion-related knowledge.
-Recent efforts have focused on improving MLLMs by incorporating multimodal emotion datasets and emotion reasoning tasks [39, 40].
-However, these models still face challenges in leveraging contextual information effectively, particularly in sentiment analysis, where understanding the emotional nuances is crucial.
-II-B Multitask Learning (MTL)
-Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others [41].
+These models are typically trained for general-purpose tasks using a standard framework that includes modal encoders, a large language model (LLM), and a mapping layer (e.g., Linear Chen et al.
+(2023), MLP Liu et al.
+(2024); Chen et al.
+(2024a), or Q-Former Li et al.
+(2023b)).
+While they excel in general tasks, MLLMs struggle with MSA and MER due to insufficient training on sentiment-specific datasets and emotion-related knowledge.
+Recent efforts have focused on improving MLLMs by incorporating multimodal emotion datasets and emotion reasoning tasks Cheng et al.
+(2024a); Lian et al.
+(2024).
+However, these models still face challenges in leveraging contextual information effectively, particularly in sentiment analysis, where understanding the nuances is crucial.
+2.2 Multitask Learning
+Multitask Learning (MTL) is a machine learning paradigm that enables models to simultaneously learn multiple related tasks, thereby allowing knowledge from one task to be transferred to others.
 In affective computing, where multimodal inputs (e.g., text, audio, and visual frames) carry diverse and distinct information, MTL frameworks aim to leverage the synergies between tasks for improved performance.
-For instance, the MTMM-ES framework [42] jointly performs sentiment and emotion analysis, showing improved results over single-task models.
-UniMSE [11] unifies MSA and ER within a single framework.
+For instance, UniMSE Hu et al.
+(2022b) unifies MSA and MER within a single framework.
 However, it relies solely on labeled data to establish the relationship between sentiment and emotion, without fully utilizing multimodal connections.
-M2Seq2Seq [43] proposes a multimodal multitask learning model based on the encoder-decoder architecture.
-Recently, Emotion-LLaMA [39] has shown promising results in MER by incorporating multitask learning, including emotion reason inference.
+M2Seq2Seq Zhang et al.
+(2023) proposes a multimodal multitask learning model based on the encoder-decoder architecture.
+Recently, Emotion-LLaMA Cheng et al.
+(2024a) has shown promising results in MER by incorporating multitask learning, including emotion cause inference.
 However, despite its success in MER, Emotion-LLaMA still struggles with MSA tasks and fails to effectively synthesize contextual information for emotion inference.
-Figure 3: Overview of the EmoVerse training process and model architecture.
-For different training stages, we assign tasks to the data based on their respective sampling rates, which are indicated in parentheses.
+Figure 3: Overview of the M2SE training strategy and EmoVerse architecture.
+For different training stages, we assign tasks to the data based on their respective sampling rates.
 The EmoVerse model is then trained using the assigned data.
 The model architecture include visual encoder, linear projector, and LLM.
-III Methodology
-This section presents our work, which consists of three key components: the construction of the EMT dataset, the formulation of the M2SE strategy, and the development of the EmoVerse model.
-III-A EMT datasets
-The EMT dataset is constructed by integrating data from the CMU-MOSEI, MELD, and ECF2.0[33] datasets to support five distinct tasks: MSA, ER, FER, ERI, and ECPC.
-Figure 2 illustrates a sample from the EMT dataset.
-For the MSA and ER tasks, we use the original sentiment labels and emotion annotations from the source datasets as task labels.
-Due to the complexity of annotating the ECPE task, we refer to the annotation method of the ECF dataset and manually annotate a portion of the data.
-This is then combined with the existing ECF2.0 data to form the ECPE portion of the EMT dataset for further emotional causality analysis.
-In the FER task, we adapt methods from the MERR dataset[39] by extracting peak frames based on the Action Units (AUs) calculated for each frame using the OpenFace tool.
+3 Affective Multitask Dataset
+The AMT dataset is constructed by integrating data from the CMU-MOSEI, MELD, and ECF2.0 Wang et al.
+(2024a) datasets to support five distinct tasks: MSA, MER, FER, ERI, and ECPC.
+Figure 2 illustrates a sample from the AMT dataset.
+For the MSA and MER tasks, we use the original sentiment labels and emotion annotations from the source datasets as task labels.
+Due to the complexity of annotating the ECPE task, we refer to the annotation method of the ECF2.0 dataset and manually annotate a portion of the data.
+This is then combined with the existing ECF2.0 data to form the ECPE portion of the AMT dataset for further emotional causality analysis.
+In the FER task, we adapt methods from the MERR dataset Cheng et al.
+(2024a) by extracting peak frames based on the Action Units (AUs) calculated for each frame using the OpenFace tool.
 The composite score for each frame is determined by summing the AU values, as shown in the following equation:
@@
 The peak frame with the highest score is selected to represent the most expressive moment.
-For videos featuring multiple characters, such as those in the MELD dataset, we improve the method by extracting peak frames separately for each character and then selecting the final peak frame by comparing the highest composite scores from all characters:
+For videos featuring multiple characters, such as MELD dataset, we improve the method by extracting peak frames separately for each character and then selecting the final peak frame by comparing the highest composite scores from all characters:
 $S_{f_{\text{final}}}=\max\left(\{S_{f}^{(1)},S_{f}^{(2)},\dots,S_{f}^{(k)}\}\right)$ (2)
@@
 Finally, we use the LLaMA 3.1-8B  [^2] https://github.com/meta-llama/llama3 model to infer the underlying causes of the characters’ emotions by synthesizing the character and scene descriptions.
-We perform a thorough preprocessing and quality control process on the raw data from the three datasets, systematically eliminating problematic or noisy data.
-Simultaneously, we conduct a manual curation of high-quality samples, following the aforementioned methodology.
-This involves refining the dataset to ensure accuracy and consistency, while extracting the necessary labels for FER and ERI.
-The final dataset, after this meticulous refinement, is outlined in Table I, including the number of tasks per category.
-Note that each data sample can be associated with multiple tasks.
-Table I below breaks down the number of tasks that can be used.
+We conduct a comprehensive preprocessing and quality control procedure on the raw data from the three datasets, systematically removing problematic or noisy samples.
+In parallel, we manually curate high-quality samples according to the methodology described above.
+The final refined dataset, detailed in Table 1, includes the number of tasks per category.
+It is important to note that each data sample may be associated with a varying number of tasks.
+Table 1 provides a detailed breakdown of the tasks available for each category.
 Additionally, we transform the dataset into a structured format, which includes three key elements: query, response, and images/videos. This format facilitates the instruction tuning of MLLMs.
-TABLE I: The Number of Entries in EMT Dataset
-Tasks Number of Entries Sentiment Analysis 25,859 Emotion Recognition 25,859 Facial Expression Recognition 15,870 Emotion Reason Inference 4,839 Emotion Cause-Pair Extraction 7,081 Total (Summed) 32,940
-III-B EmoVerse
+Tasks Number of Entries
+Multimodal Sentiment Analysis 25859
+Multimodal Emotion Recognition 25859
+Facial Expression Recognition 15870
+Emotion Reason Inference 4839
+Emotion Cause-Pair Extraction 7081
+Total 32940
+Table 1: The Number of Entrie in the AMT Dataset
+4 Exploring the MLLM design space: How do different tasks influence the performance
+In this section, we explore how different tasks and varying sampling ratios for each task impact the final performance of the MLLM.
+After identifying the optimal training strategy, we introduce EmoVerse.
+Generally, the training process for MLLMs is divided into two stages: pretraining and instruction tuning.
+Inspired by the human learning process, which progresses from simple to complex, we also divide the affective domain fine-tuning process into these two stages, adopting a training approach that goes from simple to difficult.
+We conduct experiments using EmoVerse-4B as well as the AMT dataset (refer to Section 5.2).
+In the AMT dataset, the simplest task is MSA, which requires the model to perform binary classification.
+The more challenging tasks are MER and FER, while ERI and ECPE, which involve reasoning and inference, are the most difficult.
+Based on this, our approach is to first train the model using MSA, MER, and FER in a multitask learning setup, with MSA receiving the highest sampling ratio due to its simplicity.
+In the second stage, we focus on training the model with MER, ERI, and ECPE in a multitask learning setup, building on the model’s performance from the previous stage.
+4.1 Multitask Pretraining
+In this phase, we randomly select 15,000 samples from the AMT dataset.
+The primary objective is to enable the model to recognize facial expressions and, through these expressions, capture the relationship between sentiment and emotion.
+As discussed earlier, MSA is the simplest task, and therefore, it is assigned the highest sampling ratio, while MER, being a more complex task, receives a lower sampling ratio.
+In this stage, we also investigate the impact of varying the sampling ratio for the FER task on the model’s performance.
+To assess the model’s performance during this phase, we use the CMU-MOSEI test set.
+Model Acc2 EmoVerse (6:1:3) 77.09 / 87.47 EmoVerse (6:3:1) 82.61 / 87.50
+Table 2: Effect of the FER Task. MSA:FER:MER
+As shown in Table 2, increasing the sampling rate for the FER task leads to improved performance on the MSA task.
+Consequently, for the first stage, we implement a multitask training setup with a sampling ratio of MSA:FER:MER = 6:3:1.
+This configuration enables the model to not only understand the relationship between sentiment and emotion but also accurately recognize facial action units (AUs), thereby enhancing MSA performance.
+4.2 Multitask Reason Fine-Tuning
+For this phase, we randomly select the remaining samples from the AMT dataset and distribute them across the tasks.
+As previously mentioned, in the second phase, we introduce more challenging tasks, including MER, ERI, and ECPE.
+The final performance of the MSA task is evaluated on the CMU-MOSEI dataset, the MER task performance on the MELD dataset, and the ECPE task performance on the ECF2.0 dataset (refer to Section 5.1).
+We begin with the MER task and gradually introduce additional tasks to examine the impact of training with different tasks.
+The results are summarized in Tabel 3.
+Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse (First stage) 82.61 / 87.50 - - - - EmoVerse (MER) 81.68 / 87.20 66.74 66.24 25.40 26.69 EmoVerse (MER,ERI,EPCE) 79.00 / 86.78 66.97 65.67 69.85 69.95 EmoVerse (MER,ERI,EPCE,MSA) 85.51 / 87.80 67.78 66.74 72.30 72.21
+Table 3: Effect of the Distribution of Tasks on EmoVerse.
+When we train the model in the second phase using only these three tasks, the final model performs poorly on the MSA task.
+This indicates the presence of catastrophic forgetting in MLLM multitask training.
+However, when we allocate a small portion of the MSA task in the second phase, overall performance improves.
+This suggests that revisiting MSA-related knowledge helps the model deepen its understanding of the relationship between sentiment and emotion.
+In the second phase, the final sampling ratio we chose is MER:ERI:ECPE:MSA = 3:3:3:1.
+Therefore, the optimal training strategy we have identified involves using the MSA, FER, and MER tasks in the first phase, and the MSA, MER, ERI, and ECPE tasks in the second phase for multi-stage multitask training.
+We refer to this strategy as M2SE.
+4.3 EmoVerse
 To validate the effectiveness of the M2SE strategy, we select the most widely used MLLM framework without introducing any multitask modifications to its architecture.
@@
 The proposed MLLM architecture, EmoVerse, as shown in Figure 3, consists of a visual encoder, a linear projector, and the LLM.
-We train EmoVerse following M2SE strategy using the EMT datasets.
+We train EmoVerse following M2SE strategy using the AMT datasets.
 For a data sample $D$, the input consists of the pair $\{v,t\}$, where $v$ represents the video and $t$ represents the text.
-First, the video $v$ is processed to extract visual tokens using a Vision Transformer (ViT)[44], producing a set of visual tokens $T_{v}$.
+First, the video $v$ is processed to extract visual tokens using a Vision Transformer (ViT)Dosovitskiy et al.
+(2021), producing a set of visual tokens $T_{v}$.
 These visual tokens $T_{v}$ are then passed through a linear transformation $\mathbf{W}_{v}$ to align the visual features with the feature space of the LLM.
@@
 where $\mathcal{M}$ is the model and $\hat{y}$ is the model’s prediction or answer.
-III-C Multi-Stage Multitask Sentiment and Emotion Instruction Tuning
-III-C1 Stage 1: Multitask Pretraining
-In the first phase of training, the model focuses on three key tasks: MSA, ER, and FER.
-We randomly select 15,000 samples from the EMT dataset for this phase, giving higher sampling rates to the MSA and FER tasks, while assigning a lower sampling rate to the ER task.
-The primary objective of this stage is to enable the model to recognize facial expressions and, through these expressions, capture the relationship between sentiment and emotion.
-Specifically, the FER task helps the model learn fine-grained facial features, which in turn improves performance in both the MSA and ER tasks by reinforcing the connection between emotional cues and sentiment.
-III-C2 Stage 2: Multitask Reason Fine-Tuning
-Following the first phase, where the model learns the fundamental relationship between sentiment and emotion as well as facial expression recognition, the second phase introduces four primary tasks: MSA, ER, ERI, and ECPE.
-For this phase, we select the remaining samples from the EMT dataset, randomly distributing them across the tasks.
-Notably, the MSA task is assigned a very small sampling rate, while the other three tasks—ER, ERI, and ECPE—are given higher and equal sampling rates.
-The ERI task enhances the model’s ability to reason about emotional causes by analyzing speech, facial expressions, and contextual information, building on the facial expression recognition learned in Stage 1.
-On the other hand, the ECPE task provides deeper insights into the causes of emotions by linking emotional responses with their underlying causes in conversations.
-Through the extraction of emotion cause-pairs, ECPE fosters better integration of multimodal and contextual information, helping the model understand the complex relationships between sentiment, emotion, and context.
-The details of Stage 1 and Stage 2 can be found in the Table II.
-TABLE II: Stage 1 and Stage 2 with Tasks and Task Identifiers
-Different Tasks Stages Task Identifier Stage 1 Stage 2 Sentiment Analysis <sentiment> Emotion Recognition <emotion> Facial Expression Recognition × <caption> Emotion Reason Inference × <reason> Emotion Cause-Pair Extraction × <emotion cause-pair>
-IV Experiments
-IV-A Experiments Setup
-EmoVerse is trained solely on the EMT dataset without the addition of any external data, and it is tested on the following task-specific test sets.
-IV-A1 Sentiment Analysis.
+Method CMU-MOSEI MELD ECF2.0
+Acc2(N/N) Acc2(N/P) Acc WF1 F1 WF1
+MulT - 82.50 - - - -
+Self-MM 82.81 85.17 - - - -
+MMIM 82.24 85.97 - - - -
+BBFN - 86.20 - - - -
+FDMER - 86.10 - - - -
+MISA 83.60 85.50 - - - -
+ConKI 82.73 86.25 - - - -
+ConFEDE 81.65 85.82 - - - -
+PMR - 83.30 - - - -
+MRC-D3AE 83.10 85.50 - - - -
+DialogueGCN - - 59.46 58.10 - -
+DialogueCRN - - 60.73 58.39 - -
+DAG-ERC - - - 63.65 - -
+MM-DFN - - - 59.46 - -
+EmoCaps - - - 64.00 - -
+DisGCN - - - 64.22 - -
+MMGCN - - 62.49 58.65 - -
+GA2MIF - - - 58.94 - -
+UniMSE - - 65.09 65.51 - -
+FacialMMT - - - 66.58 - -
+LLaVA-Video - 82.82 23.87 22.65 22.59 22.78
+InternVL2-4B - 82.09 39.46 40.59 33.24 33.35
+InternVL2-8B - 85.30 42.18 44.13 40.57 40.51
+EmoVerse-4B 85.51 87.80 67.28 66.58 72.30 72.21
+EmoVerse-8B 85.93 88.51 67.78 66.74 73.54 73.62
+Table 4: Comparison of EmoVerse with other SOTA methods across different tasks.
+5 Experiments
+5.1 Experiments Setup
+EmoVerse is trained solely on the AMT dataset without the addition of any external data, and it is tested on the following task-specific test sets.
+5.1.1 Sentiment Analysis
 We use the CMU-MOSEI datasets for MSA.
@@
 The results for Acc-2 calculated based on negative/non-negative (N/N) scheme, while the results on the right are calculated according to negative/positive (N/P) definition.
-IV-A2 Emotion Recognition.
+5.1.2 Emotion Recognition
 We use the MELD datasets for ER.
@@
 We use Acc and weighted f1-sorce for evaluation.
-IV-A3 Emotion Cause-pair Extraction.
+5.1.3 Emotion Cause-pair Extraction
 We use the ECF2.0 datasets for ECPE.
@@
 We refer to previous evaluations and assess the ECF2.0 test set using the F1 Score and Weighted F1 Score.
-IV-B Implementation Details
-For the visual encoder, we utilize pre-train ViT with input images resized to $448\times 448$ pixels.
-During the fine-tuning process, the visual backbone is frozen, with the focus placed on training the linear projection layer.
-For LLM, we employ Phi-3-Mini[45] for EmoVerse-4B and InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLMfor EmoVerse-8B, both augmented with LoRA[46] for parameter-efficient fine-tuning.
-In each stage of training, we train for 2 epochs with LoRA hyperparameters set to $r=8$ and $a=32$.
-The learning rate is set to $1\times 10^{-5}$, and we use a warm-up cosine schedule for the learning rate decay.
-For EmoVerse-4B, the training is conducted on 2 NVIDIA RTX 3090 GPUs, requiring approximately 48 hours.
-For EmoVerse-8B, the training is executed on a NVIDIA A100 GPU, taking around 55 hours.
-IV-C Results
-IV-C1 Performance comparison.
-For MSA task, we compare our EmoVerse model with several SOTA baselines, including TFN[7], MulT[12], PMR[21], MISA[17], MMIM[14], Self-MM [13], BBFM[15], FDMER[16],CONKI[18], DMD[20], ConFEDE[19], and MRC-D3AE[27].
-In ER task, we evaluate our model against baseline models with DialogueGCN [22], DialogueCRN[23], DAG-ERC[8], MM-DFN[24], EmoCaps[9], ERMC-DisGCN[25], MMGCN[26], GA2MIF[10], UniMSE[11], and FacialMMT[47].
-Additionally, for both MSA and ER, we select advanced models capable of processing both images and videos.
-These include LLaVA-Video [48] and InternVL2 [49], both of which are designed to handle multimodal input efficiently.
-The performance comparison results, presented in Table III and Table IV, clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, surpassing all the aforementioned baseline models in terms of accuracy and robustness.
-For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task. The Table V shows the results of ECPE.
-TABLE III: The performance comparison on CMU-MOSEI dataset.
-Model Acc2 TFN - / 82.5 MulT - / 82.5 Self-MM 82.81 / 85.17 MMIM 82.24 / 85.97 BBFN - / 86.20 FDMER - / 86.10 MISA 83.60 / 85.50 ConKI 82.73 / 86.25 ConFEDE 81.65 / 85.82 DMD - / 86.60 PMR - / 83.30 MRC-D3AE 83.10 / 85.50 LLaVA-Video - / 82.82 InternVL2-4B - / 82.09 InternVL2-8B - / 85.30 EmoVerse-4B 85.51 / 87.80 EmoVerse-8B 85.93 / 88.51
-TABLE IV: The performance comparison on MELD dataset.
-Model Acc W.F1-Score DialogueGCN 59.46 58.10 DialogueCRN 60.73 58.39 DAG-ERC - 63.65 MM-DFN - 59.46 EmoCaps - 64.00 ERMC-DisGCN - 64.22 MMGCN 62.49 58.65 GA2MIF - 58.94 UniMSE 65.09 65.51 FacialMMT - 66.58 LLaVA-Video 23.87 22.65 InternVL2-4B 39.46 40.59 InternVL2-8B 42.18 44.13 EmoVerse-4B 67.28 66.58 EmoVerse-8B 67.78 66.74
-TABLE V: The performance comparison on ECF2 dataset.
-Model W.F1-Score F1-Score LLaVA-Video 22.59 22.78 InternVL2-4B 33.24 33.35 InternVL2-8B 40.57 40.51 EmoVerse-4B 72.30 72.21 EmoVerse-8B 73.54 73.62
-IV-C2 Analysis of Emotion Reasoning.
-To demonstrate the qualitative performance of EmoVerse, we present a detailed comparison of emotion reasoning results across various models.
-Figure 4 showcases the emotion reasoning outputs of five models: GPT-4o, Video-LLaMA2[50], LLaMA-Video, InternVL2, and EmoVerse.
-The video depicts a woman displaying a strong reaction of surprise, highlighting the models’ ability to infer emotional states from multimodal inputs.
-GPT-4o and Video-LLaMA2 both classify the emotion in the video as ”confident” or ”assertive.” However, LLaMA-Video and InternVL2 correctly identify the emotion as ”surprise.” Despite this, these two models fail to fully integrate multimodal information for ERI.
-For instance, LLaMA-Video makes its inference solely based on facial expressions, while InternVL2 briefly acknowledges the facial expressions without providing a detailed explanation of the emotional cause, instead relying solely on textual cues for emotion reasoning.
-In contrast, EmoVerse effectively combines both multimodal video and text information to infer the cause of the emotion, yielding a correct and comprehensive emotional interpretation.
 Figure 4: Comparison of different models on ERI.
-It can be seen that both GPT-4o and Video-LLaMA2 make errors in ER.
+It can be seen that both GPT-4o and Video-LLaMA2 make errors in MER.
 While Llava-Video and InternVL2 correctly classify the emotions, their descriptions of the reasons behind the emotions are incomplete.
 In contrast, EmoVerse not only correctly classifies the emotions but also provides a comprehensive description of the reasons behind them.
-IV-D Ablation Evaluation
-We use EmoVerse-4B to do the following ablation experiments.
-IV-D1 Comparison of the M2SE Strategy with Single-Task Fine-Tuning.
-To evaluate the impact of the M2SE strategy within the same model framework and compare its performance across various tasks, we conduct an ablation study.
-As shown in Table VI, the tasks in parentheses represent the fine-tuning tasks, with each task corresponding to the same datasets used in our evaluation.
-”Z.S.” stands for zero-shot.
-The M2SE strategy is trained using the EMT dataset.
-The results demonstrate that, although the performance on some metrics may not exceed that of single-task learning, EmoVerse trained with the M2SE strategy consistently performs better overall compared to other models.
-TABLE VI: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
-Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Z.S.) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(ER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.28 66.58 72.30 72.21
-IV-D2 Effect of the FER Task.
-To explore the contribution of the FER task to MSA, we conduct an ablation experiment to compare the effects of increasing and decreasing the FER sampling ratio on MSA performance, using the same training data volume in the first stage.
-For a low sampling ratio, we denote it as ”lf,” and for a high sampling ratio, we denote it as ”hf.” As shown in Table VII, under the same data volume, increasing the FER task ratio—despite the model performing fewer MSA tasks—still leads to stronger learning performance.
-This indicates that even with fewer MSA samples, the contribution of FER significantly enhances the model’s ability to perform sentiment analysis.
-TABLE VII: Effect of the FER Task.
-Model Acc2 EmoVerse (lf) 77.09 / 87.47 EmoVerse (hf) 82.61 / 87.50
-IV-D3 Effect of the Choice of Tasks in the Second Stage.
-To investigate how task configurations impact model performance and better understand cross-task interactions, we conduct an ablation study in the second stage with three setups: ”t1” (only ER task), ”t3” (ER, ERI, and ECPE tasks), and ”t4” (all four tasks).
-Results, as shown in Table VIII, show that focusing solely on ER (t1) improves ER performance but weakens MSA and emotion reasoning tasks.
-Adding ERI and ECPE (t3) slightly lowers MSA and ER performance but enhances the model’s emotion reasoning ability.
-When MSA is reintroduced with a small sampling rate (t4), the model effectively integrates cross-task knowledge, improving performance across all tasks and surpassing first-stage MSA results.
-This demonstrates the importance of balanced task integration for holistic model improvement.
-TABLE VIII: Effect of the Distribution of Tasks on EmoVerse.
-Model CMU-MOSEI MELD ECF2.0
-Acc2 Acc W.F1 F1 W.F1
-EmoVerse (f) 82.61 / 87.50 - - - -
-EmoVerse (t1) 81.68 / 87.20 66.74 66.24 25.40 26.69
-EmoVerse (t3) 79.0 / 86.78 66.97 65.67 69.85 69.95
-EmoVerse (t4) 85.51 / 87.80 67.78 66.74 72.30 72.21
-V Conclusion
-In this paper, we propose the M2SE strategy to guide the training process, enabling the model to learn from multiple sentiment- and emotion-related tasks across different stages.
-To support this, we construct the EMT dataset, which includes a large number of annotated samples spanning a wide range of tasks.
-Additionally, we introduce EmoVerse, a unified MLLM designed to enhance emotion perception by effectively integrating both visual and textual data.
-By leveraging multitask and multistage learning, EmoVerse unifies sentiment and emotion tasks, achieving superior performance in MSA, ER, and EPCE.
-It outperforms existing models on benchmark datasets such as CMU-MOSEI and MELD.
-The experimental results highlight the effectiveness of the M2SE strategy in enabling the model to capture and reason about complex emotional contexts, showcasing its potential to drive advancements in emotion-aware AI applications.
+5.2 Implementation Details
+For the visual encoder, we use a pre-trained ViT with input images resized to $448\times 448$ pixels.
+During fine-tuning, the visual backbone is frozen, and training focuses on the linear projection layer.
+For the LLM, EmoVerse-4B uses Phi-3-Mini Abdin et al.
+(2024), and EmoVerse-8B uses InternLM-2.5-7B-Chat [^3] https://github.com/InternLM/InternLM, both augmented with LoRA Hu et al.
+(2021b) for parameter-efficient fine-tuning.
+Training is done for 2 epochs with LoRA hyperparameters $r=8$ and $a=32$, a learning rate of $1\times 10^{-5}$, and a warm-up cosine decay schedule.
+EmoVerse-4B training takes approximately 48 hours on 2 NVIDIA RTX 3090 GPUs, while EmoVerse-8B requires about 55 hours on a single NVIDIA A100 GPU.
+5.3 Results
+5.3.1 Performance comparison
+For MSA task, we compare our EmoVerse model with several SOTA baselines, including MulT Tsai et al.
+(2019), PMR Lv et al.
+(2021), MISA Hazarika et al.
+(2020), MMIM Han et al.
+(2021b), Self-MM Yu et al.
+(2021), BBFM Han et al.
+(2021a), CONKI Yu et al.
+(2023), ConFEDE Yang et al.
+(2023), and MRC-D3AE Zhang et al.
+(2024a).
+In MER task, we evaluate our model against baseline models with DialogueGCN Ghosal et al.
+(2019), DialogueCRN Hu et al.
+(2021a), DAG-ERC Shen et al.
+(2021), MM-DFN Hu et al.
+(2022a), EmoCaps Li et al.
+(2022), DisGCN Sun et al.
+(2021), MMGCN Hu et al.
+(2021c), GA2MIF Li et al.
+(2023a), UniMSE Hu et al.
+(2022b), and FacialMMT Zheng et al.
+(2023).
+Additionally, for both MSA and MER, we select models capable of processing both images and videos, including LLaVA-VideoZhang et al.
+(2024b) and InternVL2Chen et al.
+(2024b).
+For the ECPE task, we conduct tests solely on MLLMs, as they are better equipped to handle the complexity and multimodal nature of the task.
+The performance comparison results, shown in Table 4 clearly demonstrate that EmoVerse achieves SOTA performance across both tasks, outperforming all the aforementioned baselines.
+5.3.2 Analysis of Emotion Reasoning
+To demonstrate EmoVerse’s qualitative performance, we compare emotion reasoning results across five models: GPT-4o, Video-LLaMA2Cheng et al.
+(2024b), LLaMA-Video, InternVL2, and EmoVerse, as shown in Figure 4.
+The video depicts a woman displaying a strong surprise reaction, highlighting the models’ ability to infer emotional states from multimodal inputs.
+GPT-4o and Video-LLaMA2 misclassify the emotion as ”confident” or ”assertive,” while LLaMA-Video and InternVL2 correctly identify it as ”surprise.” However, these models fail to fully integrate multimodal information for emotion reasoning.
+LLaMA-Video relies solely on facial expressions, while InternVL2 combines facial cues with textual analysis but lacks detailed explanation of the emotional cause.
+In contrast, EmoVerse effectively combines video and text to accurately infer the emotional cause, providing a comprehensive interpretation.
+Model CMU-MOSEI MELD ECF2.0 Acc2 Acc W.F1 F1 W.F1 EmoVerse(Zero-shot) 82.09 / 83.87 39.46 40.59 33.24 33.35 EmoVerse(MSA) 83.96 / 88.16 48.12 31.27 41.67 41.14 EmoVerse(MER) 78.99 / 80.17 66.21 65.71 40.61 38.23 EmoVerse(EPCE) 83.27 / 85.85 32.38 33.43 72.69 72.59 EmoVerse(M2SE) 85.51 / 87.80 67.78 66.74 72.30 72.21
+Table 5: Comparison of the M2SE Strategy with Single-Task Fine-Tuning
+5.4 Comparison of the M2SE Strategy with Single-Task Fine-Tuning
+To assess the impact of the M2SE strategy within the same framework and compare its performance across tasks, we conduct an ablation study.
+Table 5 lists the fine-tuning tasks (in parentheses) with their corresponding datasets.
+The M2SE strategy uses the AMT dataset for training.
+Results show that while some metrics may fall short of single-task learning, EmoVerse with M2SE outperforms other models overall.
+6 Conclusion
+In this work, we explore the impact of different affective tasks on MLLM under a multistage and multitask training framework, identifying and validating the optimal strategy called M2SE.
+Using the M2SE strategy, we train EmoVerse, addressing the lack of a general-purpose MLLM in affective computing.
+Additionally, we construct the AMT dataset, designed to support sentiment and emotion perception training for MLLMs.
+Experimental results demonstrate that EmoVerse achieves state-of-the-art performance across multiple tasks.
+This study highlights the effectiveness of the M2SE strategy in unifying sentiment and emotion tasks, paving the way for more emotionally intelligent AI systems.
 Acknowledgments